OpenStudy (anonymous):

Find the value of:

OpenStudy (anonymous):

\[\frac{ 2\sqrt{3} }{ \sqrt{10}}\]

OpenStudy (anonymous):

in simplest form.

OpenStudy (anonymous):

bye

OpenStudy (austinl):

\(\dfrac{2\sqrt{3}}{\sqrt{10}}\times \dfrac{\sqrt{10}}{\sqrt{10}}=\dfrac{2\sqrt{3}\times\sqrt{10}}{10}=\dfrac{2\sqrt{30}}{10}=\dfrac{\sqrt{30}}{5}\)

OpenStudy (austinl):

This is because you cannot have a radical in the denominator, so you have to multiply both top and bottom by said radical.

OpenStudy (anonymous):

find the value of \[\frac{ \sqrt{6} }{ 5\sqrt{3} }\]

OpenStudy (anonymous):

in simplest form.

OpenStudy (austinl):

You do the same thing, multiply both top and bottom by \(5\sqrt{3}\), and then simplify from there.

OpenStudy (anonymous):

can u walk me through it

OpenStudy (austinl):

What is \(5\sqrt{3}\times5\sqrt{3}\)?

OpenStudy (anonymous):

idk

OpenStudy (austinl):

Do you have a calculator? If not, try www.wolframalpha.com

OpenStudy (anonymous):

525

OpenStudy (austinl):

Think of it like this, \(5\times 5=25\) \(\sqrt{3}\times\sqrt{3}=3\) \(25\times3=?\)

OpenStudy (anonymous):

75

OpenStudy (austinl):

Correct, and what is \(\sqrt{6}\times5\sqrt{3}\)

OpenStudy (anonymous):

\[15\sqrt{2}\]

OpenStudy (austinl):

Correct, so now we have \(\dfrac{15\sqrt{2}}{75}\) How would you simplify that?

OpenStudy (anonymous):

idk this website is freaking stupid

OpenStudy (austinl):

\(75\div 15=?\)

OpenStudy (anonymous):

5

OpenStudy (austinl):

Correct, so now we have \(\large{\dfrac{\sqrt{2}}{5}}\) \(\color{red}{D}\color{blue}{O}\color{green}{N}\color{purple}{E}\)

OpenStudy (calculusxy):

Yes you're right because 75 divided by 15 equals to 5.